How do you add fractions? Do you have to flip it or anything I can multiple but not subtract or add.
3/20+4/5*3/4
* means Multiply

Answer by stanbon(75887)

3/20+4/5*3/4
------------
The key to adding or subtractin fractions is
to create a common denominator for each fraction.
-----------------
Your Problem:
(3/20) + (3/5)
Common Denominator is 20.
Rewrite each fraction with the common denominator:
=(3/20) + (12/20)
Now you can add the fractions:
= 15/20
= 3/4
==========
Cheers,
Stan H.